# Entangelment Entropy on Generalized Brillouin Zone

### v1

@@ V Entanglement entropy on non-circular GBZ
Figure 5: The von Neumann (a) and $\alpha=2$ edge (b) entanglement entropy, calculated on the quasi-reciprocal lattice, are plotted as a function of $t_{1}$. Real and imaginary parts of entropy are shown in red and blue color, respectively. (c) The real part of the entanglement spectrum $\xi$ as a function of $t_{1}$. (d) The maximum imaginary part of entanglement spectrum at $t_{1}=1.45$ varies as a function of system length $L$. It decreases to a small value once $L$ becomes sufficiently large. The inset shows the imaginary of the entanglement spectrum. The topological transition occurs at $t_{1c}=1.562$ and is denoted by the black dashed line. The values of parameters used in the calculations are $\gamma=2/3$, $t_{2}=1$, $t_{3}=0.2$ and $L=800$.
@@
The phase diagram of quasi-reciprocal Hamiltonian is shown in Fig. 6. A gapless semimetal phase (Phase II) appears when $t_{3}\neq 0,$ in which the system becomes exceptional. We then consider possible scaling behaviors along the boundaries delineating this exceptional phase [43]. Since the von Neumann entropy $S_{A}$ is complex, we adopt the generalized entanglement entropy $S_{A}^{g}$, proposed in Ref. [38], to investigate the critical properties,
@@
We introduce the quasi-reciprocal Hamiltonian and lattice in real space, which is obtained by performing a regular Fourier transformation on the non-Bloch Hamiltonian. We calculate the entanglement entropy on these quasi-reciprocal lattices, and term it as the non-Bloch entanglement entropy. Since the quasi-reciprocal lattice is free of skin effect, the non-Bloch entanglement entropy exhibits well-behaved characteristics. For cases where the GBZ is circular, the non-Bloch entanglement entropy is shown to be well-defined (real and positive-definite). The non-Bloch entanglement entropy exhibits a logarithmic scaling in the critical region. It is found that the critical behavior is associated with the Fermi surface structure of the energy spectrum: each Fermi point contributes precisely 1 to the central charge $c$ of the logarithmic scaling. At the exceptional point, the central charge becomes negative due to the presence of the exceptional bound state. When the GBZ is non-circular, the non-Bloch von Neumann entropy becomes complex-valued in the topological region, due to the emergence of long-range hopping in the quasi-reciprocal lattice. Nonetheless, the non-Bloch edge entanglement entropy remains real and quantized, thus serving as a reliable topological indicator recovering the BBC. Furthermore, We computes the topological phase diagram and reveals the critical behavior along exceptional phase boundaries between topological, semimetal, and trivial phases. The central charge exhibits interesting behaviors, such as becoming precisely $-2$ at the critical point with $t_{3}=t_{2}$ or oscillating as a function of the system’s length on the topological-semimetal phase boundary. Overall, this study sheds new light on the physical meanings of the GBZ and non-Hermitian entanglement entropy, opening new possibilities for studying many-body physics based on the GBZ.
@@ ACKNOWLEDGMENT
Note added. After completing our work, we became aware of a preprint [61] whose findings partially overlap with ours.

### v2

@@ V Entanglement entropy on non-circular GBZ
Figure 5: The von Neumann (a) and $\alpha=2$ edge (b) entanglement entropy, calculated on the quasi-reciprocal lattice, are plotted as a function of $t_{1}$. Real and imaginary parts of entropy are shown in red and blue color, respectively. (c) The real part of the entanglement spectrum $\xi$ as a function of $t_{1}$. (d) The maximum imaginary part of entanglement spectrum at $t_{1}=1.45$ varies as a function of system length $L$. It decreases to a small value once $L$ becomes sufficiently large. The inset shows the imaginary of the entanglement spectrum. The topological transition occurs at $t_{1c}=1.562$ and is denoted by the black dashed line. The values of parameters used in the calculations are $\gamma=2/3$, $t_{2}=1$, $t_{3}=0.2$, and $L=800$.
@@
The phase diagram of quasi-reciprocal Hamiltonian is shown in Fig. 6. A gapless semimetal phase (Phase II) appears when $t_{3}\neq 0$ [69], in which the system becomes exceptional. We then consider possible scaling behaviors along the boundaries delineating this exceptional phase [46]. Since the von Neumann entropy $S_{A}$ is complex, we adopt the generalized entanglement entropy $S_{A}^{g}$, proposed in Ref. [41], to investigate the critical properties,
@@
We introduce the quasi-reciprocal Hamiltonian and lattice in real space, which is obtained by performing a regular Fourier transformation on the non-Bloch Hamiltonian. We calculate the entanglement entropy on these quasi-reciprocal lattices, and term it as the non-Bloch entanglement entropy. Since the quasi-reciprocal lattice is free of skin effect, the non-Bloch entanglement entropy exhibits well-behaved characteristics. For cases where the GBZ is circular, the non-Bloch entanglement entropy is shown to be well-defined (real and positive-definite). The non-Bloch entanglement entropy exhibits a logarithmic scaling in the critical region. It is found that the critical behavior is associated with the Fermi surface structure of the energy spectrum: each Fermi point contributes precisely 1 to the central charge $c$ of the logarithmic scaling. At the exceptional point, the central charge becomes negative due to the presence of the exceptional bound state. When the GBZ is non-circular, the non-Bloch von Neumann entropy becomes complex-valued in the topological region, due to the emergence of long-range hopping in the quasi-reciprocal lattice. Nonetheless, the non-Bloch edge entanglement entropy remains real and quantized, thus serving as a reliable topological indicator recovering the BBC. Furthermore, We compute the topological phase diagram and reveal the critical behavior along exceptional phase boundaries between topological, semimetal, and trivial phases. The central charge exhibits interesting behaviors, such as becoming precisely $-2$ at the critical point with $t_{3}=t_{2}$ or oscillating as a function of the system’s length on the topological-semimetal phase boundary. Overall, this study sheds new light on the physical meanings of the GBZ and non-Hermitian entanglement entropy, opening new possibilities for studying many-body physics based on the GBZ.
@@ ACKNOWLEDGMENT
Note added. After completing our work, we became aware of a recent paper [71]. Some of their findings overlap with our results.